OpenStudy (anonymous):

If s varies inversely as √d, find the percentage change in the value of s when the value of d is reduced by 75%?

OpenStudy (anonymous):

[-8 1/10 * \ge 1\]

OpenStudy (anonymous):

\[s_{1} = \frac{k}{\sqrt{d}}\]\[k = s_{1} \sqrt{d}\]\[s_{2} = \frac{k}{\sqrt{0.25d}}\]\[s_{2} = \frac{s_{1} \sqrt{d}}{\sqrt{0.25d}}\]\[s_{2} = s_{1}(\sqrt{\frac{1}{\frac{1}{4}}})\]\[s_{2} = s_{1}\sqrt{4}\]\[s_{2} = 2s_{1}\] THis means a 100% increase? @Mertsj Is what I did correct?

OpenStudy (anonymous):

@dpaInc Can you see if what I did is correct?

OpenStudy (anonymous):

looks good to me...

OpenStudy (anonymous):

Alright. Thanks!

OpenStudy (mertsj):

Since the constant is the constant, you can just assign a value of 1 to it and omit it from the calculation. \[\frac{s _{2}}{s _{1}}=\frac{\frac{1}{\sqrt{.25d}}}{\frac{1}{\sqrt{d}}}=\frac{\sqrt{d}}{\sqrt{.25d}}=\frac{2}{1}\]

OpenStudy (mertsj):

Sorry I got tied up on the phone.

OpenStudy (anonymous):

Alright. Thanks :)

OpenStudy (mertsj):

yw